Income statements and balance sheets data for Virtual Gaming Systems are provided below.

VIRTUAL GAMING SYSTEMS
Income Statements
For the year ended December 31
   2019    2018
  Net sales $3,555,000 $3,081,000
  Cost of goods sold 2,489,000 1,959,000
       Gross profit 1,066,000 1,122,000
  Expenses:
      Operating expenses 964,000 867,000
      Depreciation expense 39,000 31,500
      Loss on sale of land 0 8,900
      Interest expense 22,500 19,500
      Income tax expense 8,900 52,500
        Total expenses 1,034,400 979,400
  Net income $    31,600 $   142,600


VIRTUAL GAMING SYSTEMS
Balance Sheets
December 31
   2019    2018    2017
  Assets
  Current assets:
      Cash $ 214,500    $195,000    $153,000   
      Accounts receivable 88,500    90,000    69,000   
      Inventory 138,500    114,000    144,000   
      Prepaid rent 14,900    12,900    7,080   
  Long-term assets:
        Investment in bonds 114,000    114,000    0   
        Land 309,000    219,000    249,000   
        Equipment 309,000    279,000    219,000   
        Less: Accumulated depreciation (121,500) (82,500) (51,000)
          Total assets $1,066,900    $941,400    $790,080   
  Liabilities and Stockholders' Equity
  Current liabilities:
      Accounts payable $    118,200    $ 75,000    $134,780   
      Interest payable 11,700    7,800    3,900   
      Income tax payable 12,900    19,500    14,900   
Long-term liabilities:
      Notes payable 490,000    294,000    234,000   
  Stockholders' equity:
      Common stock 309,000    309,000    309,000   
      Retained earnings 125,100    236,100    93,500   
         Total liabilities and stockholders’ equity $1,066,900    $941,400    $790,080   

Required:

1. Calculate the following risk ratios for 2018 and 2019: (Round your answers to 1 decimal place.)

2. Calculate the following profitability ratios for 2018 and 2019: (Round your answers to 1 decimal place.)
Answer #1

Common Risk ratios

1) Debt to capital ratio

( Current liabilities + Long term liabilities ) / ( Total debt + Equity )

  Debt to capital ratio for 2018 ,   

( 102300 + 294000 ) / 941400

= 0.42

  Debt to capital ratio for 2019 ,

(142800 + 490000 ) / 1066900

= 0.59

2) Debt to equity ratio

Total Debt / Equity

for 2018 396300 / 545100 = 0.72

for 2019 632800/ 434100 = 1.4

3) Interest coverage ratio

EBIT / Interest expense

for 2018 286500 / 19500 = 14.6

for 2019 141000 / 22500 = 6.2

4) Financial leverage ratio

Assets / Equity

for 2018 941400/ 545100 = 1.7

for 2019 941400 / 434100 = 2.1

5) Debt to asset ratio

Total debt / Asset

for 2018 396300 / 941400 = 0.42

for 2019 632800 / 941400 = 0.67

Profitability ratios

1) Gross profit margin = Gross profit / sales

For  2018 1122000 / 3081000 = .36

For 2019 1066000 / 3555000 = .29

2) Operating profit = Operating income / Sales

For  2018 255000 / 3081000 = .08

For 2019 102000/ 3555000 = .02

3) Net profit margin = Net income / Sales

For  2018 142600/ 3081000 = .04

For 2019 31600 / 3555000= .0088